# Black hole X-ray binary A062000 in quiescence: hints of Faraday rotation of near-infrared and optical polarization?

### v1

@@ 4.1 The source of the polarization
Next, we consider the scenario where the observed polarization arises from two components: one coming from the scattering of the stellar light and the other from the polarized contribution of the additional UV component. Stellar polarization is assumed to have constant PD and PA which implies a constant fraction of the scattered radiation, and spectrum as that of the optical companion (blue solid line in Fig. 9a). The additional UV component (dashed orange line), that was used to describe the UV part of the observed SED, is set to have constant PD and PA, whose values were free to vary. The blue lines in Fig. 9b,c show the dependence of PD and PA for this two-component scenario. The PD dependence on energy is noteworthy: appears higher than and due to the narrow absorption of the stellar spectrum (and its scattering) in this band. We find the parameters , for the stellar scattering and with for the UV component reproduce the energy dependence of PD and PA quite well (see Fig. 9b,c), especially in the optical band, where two components have comparable contribution to the observed flux. The model is also generally consistent with the NIR points (taken decades earlier); some deviation in the band might also arise from the Faraday effect or from the different treatment of the interstellar contribution to polarization.
The polarization of the UV component may arise from the first (single) Compton up-scattering of the disk or synchrotron photons by electrons in the hot accretion flow, similar to the case of MAXI J1820+070 (Poutanen et al., 2022). The difference of PA in the scattered and UV components can then indicate misalignment by (or ) between the axes of these components. This scenario predicts rise of PD and further rotation of PA towards the UV band (see Fig. 9). It can be tested in the future, by measuring polarization of A062000 simultaneously from NIR up to UV.
@@ 4.2 The source of the polarization variations
In addition to the changes in the polarization between different spectral states, we also see the orbital variations of the polarization. Although the orbital modulation of the polarization observed by Dolan & Tapia (1989) has not been confirmed (Dubus et al., 2008), we see the variations at the level of at the timescales of hours (see Fig. 5). To confidently claim that the polarization of A062000 is variable with the orbital phase one needs to cover at least two consecutive orbital periods with the high-precision polarimetric observations (which is a rather difficult task given the brightness of the object in quiescence and its short orbital period), we can still cautiously state that the observed polarization variability is related to the orbital phase rather than having a stochastic nature. Indeed, the polarization of A062000, folded with the orbital phase shows the pattern, typical for binary stars – double-sinusoidal variations with the PD reaching minima in the conjunctions (phases 0 and 0.5) and maxima in the quadratures (phases 0.25 and 0.75), arising from Thomson scattering of the stellar radiation on the matter, following the compact object in its orbital motion (Brown et al., 1978; Kravtsov et al., 2020), and similar to these observed in another BH X-ray binary Cyg X-1 (Kemp et al., 1979; Kravtsov et al., 2023b). In Fig. 10, we see fairly good agreement of the PD variations with the expected behavior (blue line shows the model of a scattering cloud on a circular orbit, see Appendix of Kravtsov et al. 2020). In addition, our first and second observations, performed during two different orbits are consistent with each other (see Fig. 5).
@@
The optical light curve shown in Fig. 10 has an asymmetric profile – the peak and the dip at the first half of the orbit are brighter than the other peak and dip. The so-called ellipsoidal variations of the flux, produced by the tidally distorted optical star (shown in blue dashed line at Fig. 10) cannot produce such asymmetry, therefore an additional source of flux variations is needed. There are two alternatives: either there is an additional bright component visible only at the first part of the orbit and generating additional flux (e.g. phase-dependent disk/bright spot, see e.g. Haswell et al. 1993), or some object blocks the light of the optical companion, reducing the total flux at phases from 0.5 to 1.0 (e.g. dark spots at the surface of the star, e.g. Cherepashchuk et al. 2019). We see similar asymmetry in the orbital polarization profile of A062000 – while around orbital phases from 0.5 to 1.0 the PD shows good agreement with the Thomson scattering model, at the first part of the orbit the PD is consistent with being constant. The additional bright unpolarized component, visible only at phases from 0.0 to 0.5 could reduce the polarization at these phases and explain both the asymmetric PD and flux variations, therefore we argue that the bright spot/phase-dependent disk model is more likely than the dark star spot model. Another alternative is that the asymmetric PD profile is produced by scattering of stellar radiation off slightly tilted accretion disk. In this scenario, the illuminated part of the accretion disk is visible to the observer only for half of the orbit (see Kravtsov et al., 2023b). However, scattering alone is not enough to reproduce increase of flux observed at the first half of the orbit, and the quality of our polarimetric data is not sufficient to discriminate between the complex models.
@@ 5 Conclusions
In this paper, we presented new high-precision phase-resolved optical polarimetric and quasi-simultaneous NIR-to-UV photometric observations of BH X-ray binary A062000 in a passive quiescent state. We determined and subtracted the interstellar polarization, which allowed us to derive the intrinsic polarization of A062000. Combined with the NIR polarimetric observations from Russell et al. (2016), we found that the orbit-average intrinsic PA rotates with the wavelength, changing from in the filter to in , while the PD remains at level throughout this spectral range. Folding our polarimetric observations with the orbital period, we found the significant orbital variability of polarization properties in the band. The shape of the variations with two minima and two maxima per period suggest that the polarization is most probably produced by Thomson scattering of the companion star emission off the matter which follows the BH in its orbital motion (e.g. scattering off the accretion disk/stream). The lack of variations at the first part of the orbit suggests that either we see a depolarization effect caused by the bright spot visible only at these phases or (less probable) the scattering material producing the polarization is obscured or tilted relative to the orbital axis. However, more high-precision polarimetric observations during at least several consecutive orbital periods are needed to draw unambiguous conclusions about the geometry of the scattering medium.
@@
The flat spectrum of the polarization and the presence of the orbital variations suggest that NIR-to-optical polarization has a scattering origin. However, in that case, the PA is not expected to change with the wavelength as observed. We first considered the possibility that stellar radiation scattered off the accretion disk experiences the Faraday rotation while traveling toward the observer through the magnetized plasma surrounding the accretion disk. The estimated values of magnetic field and optical depth can be consistent with realistic estimates if we consider our optical points in combination with the NIR points taken almost ten years ago. On the other hand, the steep trend of our optical points favour higher Faraday rotation rate, resulting in higher values for the line-of sight magnetic field.
@@
We considered the scenario of two polarized components having different PDs and PAs. One component is coming from the stellar scattered light and the other associated with the additional UV component seen in the spectrum. The second component may potentially arise from Compton up-scattering of the disk or synchrotron photons in the hot inner flow, similar to that in the quiescent-state low-mass X-ray binary MAXI J1820+070. The PAs of these components differ by , which translates to either or misalignment between their axes of symmetry. Future simultaneous polarimetric observations covering NIR-to-UV range would be extremely helpful in making an unambiguous choice between the models.
@@ Acknowledgements.
- Arnaud (1996) Arnaud, K. A. 1996, in ASP Conf. Ser., Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes (San Francisco: Astron. Soc. Pac.), 17
@@
- Dolan & Tapia (1989) Dolan, J. F., & Tapia, S. 1989, PASP, 101, 1135
@@
- McClintock & Remillard (1986) McClintock, J. E., & Remillard, R. A. 1986, ApJ, 308, 110
- McClintock & Remillard (2000) —. 2000, ApJ, 531, 956
@@
- Steele et al. (2004) Steele, I. A., Smith, R. J., Rees, P. C., et al. 2004, in Proc. SPIE, Vol. 5489, Ground-based Telescopes, ed. J. M. Oschmann Jr., 679

### v2

@@ 4.1 The source of the polarization
Next, we considered the scenario in which the observed polarization arises from two components: one coming from the scattering of the stellar light and the other from the polarized contribution of the additional UV component. The scattering component is assumed to have constant PD and PA, which implies a constant fraction of the scattered radiation, and spectrum as that of the optical companion (solid blue line in Fig. 9a). The additional UV component (dashed orange line), which was used to describe the UV part of the observed SED, is set to have constant PD and PA, whose values were free to vary. The blue lines in Fig. 9b and c show the dependence of PD and PA for this two-component scenario. To reproduce the rotation of the PA from the NIR to the band, the components must have different PAs: PA150° and PA190°. Because the components have comparable contribution to the optical flux, the PD of the UV component (PD1.8%) should be significantly higher than that of the scattering component (PD0.8%) to dominate the PA in . The polarization of the UV component can arise from the first (single) Compton up-scattering of the disk or synchrotron photons by electrons in the hot accretion flow, similar to the case of MAXI J1820+070 (Poutanen et al., 2022). The difference of PA in the scattered and UV components can then indicate misalignment by 50° (or 40°) between the axes of these components. This scenario predicts rise of PD and further rotation of PA toward the UV band (see Fig. 9).
While this model fairly well describes the rotation of the optical PA, it cannot reproduce the dataset in full detail. First of all, the intrinsic PAs in and differ from the model predictions. This difference may arise due to the fact that the observations in NIR were preformed a decade prior to the optical campaign while the source may have been in a different state (loop or active, Cantrell et al. 2008). There is evidence that in active quiescent state, the additional component contributes to the NIR part of the spectrum (Froning et al., 2007; Dinçer et al., 2018; Cherepashchuk et al., 2019), which could be a jet, an accretion stream-disk impact point, or a dusty circumbinary disk (Muno & Mauerhan, 2006; Gallo et al., 2007; Cherepashchuk et al., 2019). This additional NIR component, not present in our passive state observations, could however be responsible for an additional polarization in the band in the epoch of the Russell et al. (2016) observations; that would explain the observed difference in the PA. Some deviations in NIR may also arise from the Faraday rotation. Additionally, although the model describes the flat PD spectrum from NIR to reasonably well and replicates the PD rise due to the absorption line in well, the PD in is about 0.3% lower than the level predicted by the model. The reason for the apparent drop of the PD in is unknown, but we argue that the actual broadband spectrum of an object obtained simultaneously with polarization measurements should be used instead of the model spectrum of the companion star to analyze such subtle effects. Polarimetric observation in the UV band, in which we expect significant increase in the PD, can unambiguously infer the nature of PA rotation in A062000 and, subsequently, shed light on the nature of the UV excess in the binary.
4.2 On the synchrotron origin of the polarization
Synchrotron emission has been suggested to contribute to the polarization of A062000 in NIR in addition to scattering (Russell et al., 2016). By assuming the scattering component to have polarization with PA and PD and by subtracting it from the observed polarization, authors got the PD of synchrotron emission to be . Then, by assuming the jet contribution in to be 8–37%, authors found the jet to be polarized at the level of 3–18%. However, we argue that such a separation of polarized components cannot be done unambiguously with the existing data for several reasons.
First, the PA changes significantly within the optical range, meaning that polarization cannot be produced by scattering alone. Therefore, the PA of the scattered component cannot be determined explicitly with the optical data only. In previous section, we estimated the PA of scattered component to be closer to , which, if subtracted from observed polarization, would result in much smaller PD of synchrotron emission 0.8% with the PA of 180°. We note also that although the observed PA in seems to be perpendicular to the possible direction of the radio jet (see Fig. 4), after subtracting the scattering contribution, the PA of remaining synchrotron component differs from the jet axis by about , which cannot be easily explained with the jet model. Second, there is no direct evidence of the presence of NIR excess in the photometric data in passive quiescence. The apparent increase in the flux in NIR relative to blackbody is well described by the stellar spectrum alone – the “bump” at and is well pronounced in the K5V companion spectrum in Fig. 6. Therefore, it is rather difficult to determine the contribution of jet emission to the observed spectrum, especially given the strong variability of the object both within and between states. Adding the jet as a third component in addition to the scattering and UV components to the modeling is hardly justified given the quality and the number of polarimetric observations. Taking all of the above into account, although we do not rule out the possibility that jet contributes to the NIR polarization, we argue that its accurate determination is a challenging task that requires more precise simultaneous multiwavelength polarimetric observations than we have to date.
@@ 4.3 The source of the polarization variations
In addition to the changes in the polarization between different spectral states, we also see intra-day variability of the polarization. Although the orbital modulation of the polarization observed by Dolan & Tapia (1989) has not been confirmed (Dubus et al., 2008), we see the variations at a level of on timescales of hours (see Fig. 5). To confidently claim that the polarization of A062000 is variable with the orbital phase, one would need to cover at least two consecutive orbital periods with high-precision polarimetric observations (which is a rather difficult task given the brightness of the object in quiescence and its short orbital period). That said, we can cautiously state that the observed polarization variability is related to the orbital phase rather than having a stochastic nature. Indeed, the polarization of A062000, folded with the orbital phase, shows the pattern typical for binary stars – double-sinusoidal variations with the PD reaching minima in the conjunctions (phases 0 and 0.5) and maxima in the quadratures (phases 0.25 and 0.75). This is due to Thomson scattering of the stellar radiation on the matter, which follows the compact object in its orbital motion (Brown et al., 1978; Kravtsov et al., 2020), and is similar to what is observed in other BH (Kemp et al., 1979; Gliozzi et al., 1998; Kravtsov et al., 2023b) and neutron star (Egonsson & Hakala, 1991; Combi et al., 2004; Baglio et al., 2016) X-ray binaries. In Fig. 10 we see a fairly good agreement of the PD variations with the expected behavior (the blue line shows the model of a scattering cloud on a circular orbit; see the appendix of Kravtsov et al. 2020). In addition, our first and second observations, performed during two different orbits are consistent with each other (see Fig. 5). We note also that Stokes parameters of A062000, folded with the orbital period, show scatter, similar to one observed in Dubus et al. (2008) – while Stokes shows more pronounced orbital variability, Stokes seems to behave more erratically. The amplitude of this scatter, however, is comparable to the measurement errors, which complicates its interpretation, but one explanation could be the presence of inhomogeneities (clumps) in the scattering matter.
@@
The optical light curve shown in Fig. 10 has an asymmetric profile – the peak and the dip at the first half of the orbit are brighter than the other peak and dip. The so-called ellipsoidal variations of the flux, produced by the tidally distorted optical star (shown by the dashed blue line at Fig. 10), cannot produce such asymmetry; therefore, an additional source of flux variations is needed. There are two alternatives: either there is an additional bright component visible only at the first part of the orbit and generating additional flux (e.g., non-circular accretion disk or bright spot in the accretion stream; see, e.g., Haswell et al. 1993), or some object blocks the light of the optical companion, reducing the total flux at phases from 0.5 to 1.0 (e.g., dark spots at the surface of the star; e.g., Cherepashchuk et al. 2019). We see similar asymmetry in the orbital polarization profile of A062000 – while around orbital phases from 0.5 to 1.0 the PD shows good agreement with the Thomson scattering model, at the first part of the orbit the PD is consistent with being constant. The additional bright unpolarized component, visible only at phases from 0.0 to 0.5, could reduce the polarization at these phases and explain both the asymmetric PD and flux variations; therefore, we argue that the bright spot or phase-dependent disk models are more likely than the dark star spot model. Another alternative is that the asymmetric PD profile is produced by scattering of stellar radiation off slightly tilted accretion disk. In this scenario, the illuminated part of the accretion disk is visible to the observer only for half of the orbit (see Kravtsov et al., 2023b). However, scattering alone is not enough to reproduce the increase in flux observed at the first half of the orbit, and the quality of our polarimetric data is not sufficient to discriminate between the complex models.
@@ 5 Conclusions
We have presented new high-precision phase-resolved optical polarimetric and quasi-simultaneous NIR-to-UV photometric observations of BH X-ray binary A062000 in a passive quiescent state. We determined and subtracted the IS polarization, which allowed us to derive the intrinsic polarization of A062000. Using that combined with the NIR polarimetric observations from Russell et al. (2016), we found that the orbit-average intrinsic PA rotates with the wavelength, changing from in the filter to in , while the PD remains at the level throughout this spectral range. Folding our polarimetric observations with the orbital period, we found the significant orbital variability of polarization properties in the band. The shape of the variations, with two minima and two maxima per period, suggests that the polarization is most probably produced by Thomson scattering of the companion star emission off the matter, which follows the BH in its orbital motion (e.g., scattering off the accretion disk or stream). The lack of variations during the first part of the orbit suggests either that we see a depolarization effect caused by the bright spot visible only at these phases or (less probably) that the scattering material producing the polarization is obscured or tilted relative to the orbital axis. However, more high-precision polarimetric observations from at least several consecutive orbital periods are needed to draw unambiguous conclusions about the geometry of the scattering medium.
@@
The flat spectrum of the polarization and the presence of the orbital variations suggest that NIR-to-optical polarization has a scattering origin. However, in that case, the PA is not expected to change with the wavelength as observed. We first considered the possibility that stellar radiation scattered off the accretion disk experiences Faraday rotation while traveling toward the observer through the magnetized plasma surrounding the accretion disk. The estimated values of the magnetic field and optical depth can be considered consistent with realistic estimates if we take into account our optical points in combination with the NIR points taken almost ten years ago. On the other hand, the steep trend of our optical points favors a higher Faraday rotation rate, resulting in higher values for the line-of-sight magnetic field.
@@
We considered the scenario of two polarized components that have different PDs and PAs. One component is coming from the stellar scattered light and the other is associated with the additional UV component seen in the spectrum. The second component may potentially arise from Compton up-scattering of the disk or synchrotron photons in the hot inner flow, similar to what is seen in the quiescent-state low-mass BH X-ray binary MAXI J1820+070. The PAs of these components differ by 40°, which translates to either a or misalignment between their axes of symmetry. Future simultaneous polarimetric observations covering the NIR-to-UV range would be extremely helpful in unambiguously determining which of the two models best represents A062000.
@@ Acknowledgements.
- Arnaud (1996) Arnaud, K. A. 1996, in ASP Conf. Ser., Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes (San Francisco: Astron. Soc. Pac.), 17–20
- Baglio et al. (2016) Baglio, M. C., D’Avanzo, P., Campana, S., et al. 2016, A&A, 587, A102
- Baglio et al. (2014) Baglio, M. C., Mainetti, D., D’Avanzo, P., et al. 2014, A&A, 572, A99
- Bessell (1979) Bessell, M. S. 1979, PASP, 91, 589
- Boyd et al. (2001) Boyd, P. T., Smale, A. P., & Dolan, J. F. 2001, ApJ, 555, 822
@@
- Dolan & Tapia (1989) Dolan, J. F. & Tapia, S. 1989, PASP, 101, 1135
@@
- Froning & Robinson (2001) Froning, C. S. & Robinson, E. L. 2001, AJ, 121, 2212
- Froning et al. (2007) Froning, C. S., Robinson, E. L., & Bitner, M. A. 2007, ApJ, 663, 1215
- Fukugita et al. (1996) Fukugita, M., Ichikawa, T., Gunn, J. E., et al. 1996, AJ, 111, 1748
@@
- Gliozzi et al. (1998) Gliozzi, M., Bodo, G., Ghisellini, G., Scaltriti, F., & Trussoni, E. 1998, A&A, 337, L39
@@
- Koch Miramond & Naylor (1995) Koch Miramond, L. & Naylor, T. 1995, A&A, 296, 390
@@
- McClintock & Remillard (1986) McClintock, J. E. & Remillard, R. A. 1986, ApJ, 308, 110
- McClintock & Remillard (2000) McClintock, J. E. & Remillard, R. A. 2000, ApJ, 531, 956
- Muno & Mauerhan (2006) Muno, M. P. & Mauerhan, J. 2006, ApJ, 648, L135
@@
- Simmons & Stewart (1985) Simmons, J. F. L. & Stewart, B. G. 1985, A&A, 142, 100
- Steele et al. (2004) Steele, I. A., Smith, R. J., Rees, P. C., et al. 2004, in Proc. SPIE, Vol. 5489, Ground-based Telescopes, ed. J. M. Oschmann Jr., 679–692
- Tanaka et al. (2016) Tanaka, Y. T., Itoh, R., Uemura, M., et al. 2016, ApJ, 823, 35